Calculate the energy associated with the first orbit of He+what is the radius of this orbit?
1
Expert's answer
2020-12-01T09:33:39-0500

(a) calculate the energy associated with the first orbit of He+E=−13.6n2
z2​=−13.6×12
22​=54.4 evRadius of orbit r=529×z
n2​ nm=529×2
12​=264.5 nm(b) lower value of n=5s p d f g